# Lens Distortion Encoding System Version 1.0

### v1

@@ 5.1. Compositing Example
The compositing software used is Adobe After Effects with the B.Map plug-in. All assets are imported into the program, ensuring that color management is disabled for the maps. The composition is set to 32-bit floating-point precision to maintain the proper image quality.
@@
Next, we rectify the footage and track the camera motion for use in our 3D software. To begin, we create a composition using the rectilinear view map and import the footage along with its corresponding footage map. The B.Map effect is applied to the rectilinear view map layer. We flip the vertical tiling and adjust the tiling scale to account for the FOV mismatch between the view map and the footage map (see equation No. 3).
The sampled texture is then set to the footage map layer, followed by applying a second B.Map effect, which samples the footage layer. The result is a layer with rectified footage ready for camera tracking.
@@
On the pre-composed layer, the B.Map effect can be applied to sample the footage map. Following the previous steps, the warp-sampled footage can then be finalized to achieve the desired look.
@@
To enhance the quality of footage warping, 4$\times$ MSAA filtering can be applied. This process involves pre-composing the animated view map with the footage map and increasing the resolution of the pre-composition by a factor of 4. The final B.Map effect, used for sampling the footage, is then applied to this high-resolution pre-composition.

### v2

@@ 4. Mathematical Foundations
Rotating View Map Camera
View map coordinates can be rotated to simulate camera pan, tilt, and roll motion effects. This is achieved by mapping the view map coordinates to incidence vectors, which can then be rotated.
[…]
@@ 5.1. Compositing Example
The compositing software used is Adobe After Effects with the PixMap plug-in.(Wunkolo, 2024) All assets are imported into the program, ensuring that color management is disabled for the maps. The composition is set to 32-bit floating-point precision to maintain the proper image quality. The project color Working Space is configured as HDTV (Rec. 709) with enabled Linearization of Working Space. This setup ensures correct vignette blending.
@@
Next, we rectify the footage and track the camera motion for use in our 3D software. To begin, we create a composition using the rectilinear view map and import the footage along with its corresponding footage map. The PixMap effect is applied to the rectilinear view map layer. We flip the vertical tiling and adjust the tiling scale to account for the FOV mismatch between the view map and the footage map (see equation No. 3).
The sampled texture is then set to the footage map layer, followed by applying a second PixMap effect, which samples the footage layer. The result is a layer with rectified footage ready for camera tracking.
@@
On the pre-composed layer, the PixMap effect can be applied to sample the footage map. Following the previous steps, the warp-sampled footage can then be finalized to achieve the desired look.
@@
To enhance the quality of footage warping, 4$\times$ MSAA filtering can be applied. This process involves pre-composing the animated view map with the footage map and increasing the resolution of the pre-composition by a factor of 4. The final PixMap effect, used for sampling the footage, is then applied to this high-resolution pre-composition.